OpenStudy (anonymous):

Will Fan And Metal If You Help Two lines, A and B, are represented by the equations given below: Line A: y = x − 2 Line B: y = 3x + 4 Which of the following shows the solution to the system of equations and explains why? (−3, −5), because the point does not lie on any axis (−3, −5), because the point satisfies both equations (−4, −8), because the point satisfies one of the equations (−4, −8), because the point lies between the two axes

OpenStudy (anonymous):

Please i have bumped this 5 times

OpenStudy (anonymous):

I don't understand it!

Vocaloid (vocaloid):

Line A: y = x − 2 Line B: y = 3x + 4 x -2 = 3x + 4 solve for x, then find y by plugging in x into one of the equations

OpenStudy (anonymous):

i dont understand